# FIT-GNN: Faster Inference Time for GNNs that ‘FIT’ in Memory Using Coarsening

### v4

@@ 1 Introduction
To mitigate the high computational cost associated with the inference phase, we decompose the original graph $G$ into a collection of subgraphs $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$. This transformation yields substantial improvements in inference efficiency, achieving up to a $\mathbf{100\times}$ speedup for single-node prediction. Moreover, it significantly reduces memory consumption by as much as $\mathbf{100\times}$ —thereby enabling scalable inference on large-scale graphs comprising millions of edges. We provide a theoretical characterization of the upper bound on the coarsening ratio under which our method’s time and space complexity remain superior to that of the standard approach. Notably, we observe that conventional methods often fail to perform inference on large graphs due to memory constraints, whereas our approach remains effective and scalable in such scenarios.
@@ 2 Related Work
The aim of reducing computational costs has been approached through various methods. Research by 10.1145/3292500.3330925 and Zeng2020GraphSAINT: uses a subgraph sampling technique, where a small subgraph is sampled at each training iteration for GNN training. Other methods, including layer-wise sampling and mini-batch training, have been explored by chen2018stochastic; chen2018fastgcn; 10.1145/3394486.3403192; DBLP:journals/corr/abs-1911-07323. Additionally, DBLP:journals/corr/abs-2007-02817 introduced a coarsening algorithm based on Schur complements to create node embeddings in large graphs, while 10.1145/3447548.3467256 [SGGC] applied GNNs to coarsened graphs, reducing nodes and edges to lower training time. pmlr-v202-kumar23a; joly2024graph; NEURIPS2024_733209a1; 10.1145/3589335.3651920 also proposed different coarsening algorithms for scalable training approaches. jin2021graph [GCOND], wang2024fast [GC-SNTK] presented algorithms to generate a synthetic graph that mimics the original graph. However, their approach generates model-specific synthetic graphs that require training on the whole dataset, which defeats the purpose of reducing computational burden. gupta2025bonsai [BONSAI] addressed this issue by looking into computational trees and training on just the relevant computational trees. However, these approaches do not decrease inference time since the inference phase takes the whole graph as input. DBLP:journals/corr/abs-2202-00075 shifted the training process to subgraph-level, using coarsening algorithms to split the graph into $k$-subgraphs and recommending a multi-device training approach. In the graph-level task, Jin_2022 [DOSCOND] and 10.1145/3580305.3599398 [KIDD] tackled the graph classification problem following up on the approach of GCOND and GC-SNTK. However, the graph regression problem was not explored in this direction.
@@ 3 Preliminaries
3.1 Graph Notations
Given an undirected graph $G=(V,E,X,W)$, $V$ is the vertex set, $E$ is the edge set, $X\in\mathbb{R}^{n\times d}$ is the feature matrix, and $W$ is the edge weight matrix. Let $|V|=n$ be the number of nodes and $|E|=m$ be the number of edges. Let $A\in\{0,1\}^{n\times n}$ be the adjacency matrix of $G$, $d_{i}$ to be the degree of the node $v_{i}$. $D\in\mathbb{R}^{n\times n}$ is a diagonal degree matrix with $i$th diagonal element as $d_{i}$. We denote $\mathcal{N}_{j}(v_{i})$ as the $j$-hop neighbourhood of node $v_{i}\in V$.
3.2 Graph Neural Network
Graph Neural Network is a neural network designed to work for graph-structured data. The representation of each node in the graph is updated recursively by aggregation and transforming the node representations of its neighbors. kipf2017semisupervised introduced Graph Convolutional Network (GCN) as follows:
@@ 3.2 Graph Neural Network
3.3 Graph Coarsening and Graph Partitioning
loukas2018graphreductionspectralcut discussed multiple coarsening algorithms which create a coarsened graph $G^{\prime}=(V^{\prime},E^{\prime},X^{\prime},W^{\prime})$ from a given graph $G=(V,E,X,W)$. $X^{\prime}$ is the coarsened node representation. We refer the vertex set $V^{\prime}=\{v^{\prime}_{1},v^{\prime}_{2},\dots,v^{\prime}_{k}\}$ of $G^{\prime}$ as coarsened nodes. Given a coarsening ratio $r\in[0,1]$, we have $n\times r, kk\in\mathbb{Z}$. We can interpret it as creating $k$ disjoint partitions, $C_{1},C_{2},\dots,C_{k}$, from a graph of $n$ nodes. Mathematically, we create a partition matrix $P\in\{0,1\}^{n\times k}$, where $P_{ij}=1$ if and only if node $v_{i}\in C_{j}$. $A^{\prime}=P^{T}AP$ is the adjacency matrix of $G^{\prime}$. Similarly, the corresponding degree matrix $D^{\prime}=P^{T}DP$. SGGC (10.1145/3447548.3467256) used normalized partition matrix $\mathcal{P}=PC^{-\frac{1}{2}}$, where $C$ is defined as a diagonal matrix with diagonal entries $C_{jj}=|C_{j}|,j=1,2,\dots,k$.
Along with $G^{\prime}$, we create a set of disjoint subgraphs $\mathcal{G}_{s}=\{G_{1},G_{2},$ $\dots,G_{k}\}$ corresponding to the partitions $C_{1},C_{2},\dots,C_{k}$. The number of nodes in each partition $C_{i}$ is denoted by $n_{i}$. Each subgraph $G_{i}$ is the induced subgraph of $G$ formed by the nodes in $C_{i}$. $A_{i}$ and $D_{i}$ are the adjacency matrix and degree matrix of $G_{i}$ respectively.
@@ 4 Methodology
In SGGC, a graph $G$ is reduced to $G^{\prime}$ using a partition matrix $P$ generated from the coarsening algorithm. Here $X^{\prime}=P^{T}X$. The labels for the coarsened graph are $Y^{\prime}=\arg\max(P^{T}Y)$, where $Y$ is the label matrix of $G$ storing the one-hot encoding of the label of each node for the node classification task. Training is carried out on the coarsened graph $G^{\prime}$ using $Y^{\prime}$ as the label information. While this approach is novel, essential label information is lost. For instance, suppose a node $v^{\prime}\in G^{\prime}$ is a combination of $v_{i},v_{j},v_{k}\in G$ where $v_{i},v_{j},v_{k}$ are test nodes. Let us assume that the classes of $v_{i},v_{j},v_{k}$ are $0,0,2$ respectively. $\arg\max$ will take the majority label and assign class 0 to $v^{\prime}$. Hence, the model will be trained to predict only the class 0, leading to the discarding of the quantification of the model’s performance on predicting less represented nodes.
@@
  • Extra Nodes: Adding the $1$-hop neighbouring nodes, namely Extra Nodes, in each subgraph $G_{i}$ denoted by $\mathcal{E}_{G_{i}}$. DBLP:journals/corr/abs-2202-00075 introduced this approach to reduce information loss after partitioning. We add a unit weight edge if two nodes in $\mathcal{E}_{G_{i}}$ are connected in $G$. Node embedding of $v\in\mathcal{E}_{G_{i}}$ is $x_{v}\in X$.
@@
  • Cluster Nodes: Although adding $1$-hop neighboring nodes in the subgraph helps recover lost information after partition, for multi-layered GNN models, information loss will persist. Instead of adding the neighboring nodes, we will add Cluster Nodes, which represent the neighboring clusters, denoted by $\mathcal{C}_{G_{i}}$. liu2024scalable introduced this method to add nodes in the graph instead of adding to separate subgraphs. They also added cross-cluster edges among these Cluster Nodes and proposed a subgraph sampling-based method on the modified graph. In our work, we add cross-cluster edges. For each node $t\in\mathcal{C}_{G_{i}}$, the node embedding $x_{t}=X^{\prime}_{t}$.
@@
  • From Equation 2 and 3, we can say that $\sum_{i}|\mathcal{E}_{G_{i}}|\geq\sum_{i}|\mathcal{C}_{G_{i}}|$. This is because, using Extra Nodes, we add all the neighboring nodes. However, using Cluster Nodes, we only add one representative node for each neighboring cluster. Hence, the time taken to train and infer these modified subgraphs with Cluster Nodes is less than or equal to that with Extra Nodes.
@@
3:Construct $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$ using $P$;
@@
[…]
@@
8:Construct feature matrix $\{X_{1},X_{2},\dots,X_{k}\}$ for $\mathcal{G}_{s}$;
9:Construct label matrix $\{Y_{1},Y_{2},\dots,Y_{k}\}$ for $\mathcal{G}_{s}$ ;
@@
11:for each subgraph $G_{i}$ in $\mathcal{G}_{s}$ do
12:  $O_{i}=M(L,A_{i},D_{i},X_{i})$; $\quad\#$ Algorithm 4
13:  $O.append(O_{i}[mask_{i}])$;
14:  $Y.append(Y_{i}[mask_{i}])$;
15:end for
@@
17:Train $M$ to minimize LossLoss;
@@ 4.1 Node-level Task
We use Algorithm 4 with $L$ number of layers to create a standard GNN model $M$ for node-level tasks. While we use Algorithm 3 to train a model on the coarsened graph $G^{\prime}$, we propose Algorithm 1 to train a model on the set of subgraphs $\mathcal{G}_{s}$. Algorithm 3 utilizes the partition matrix $P$ generated by the coarsening algorithm to construct and train on the coarsened graph $G^{\prime}$ and the coarsened label information $Y^{\prime}$. On the other hand, our proposed Algorithm 1 uses the same partition matrix $P$ to construct and train on subgraphs $\mathcal{G}_{s}$ using original label information $Y$. For the node classification task, $\arg\max$ is used for creating labels for $G^{\prime}$ and CrossEntropy as a loss function. For the node regression task, we do not create $G^{\prime}$. Mean Absolute Error (MAE) is used as the loss function.
@@ 4.2 Graph-level Task
For a given graph $G\in\mathcal{D}$, we create $G^{\prime}$ and $\mathcal{G}_{s}=\{G_{1},G_{2},$ $\dots,G_{k}\}$. Since $G^{\prime}$ is a single graph and $\mathcal{G}_{s}$ is a set of subgraphs, we use Algorithm 5 to create a model to train on $G^{\prime}$ and Algorithm 2 to create a model to train on $\mathcal{G}_{s}$. We use the CrossEntropy loss function for the graph classification task and the Mean Absolute Error(MAE) for the graph regression task.
@@
[…]
@@
[…]
@@ 4.3 Time and Space Complexity
Train Inference (Full Graph)
@@
In Equation 1, the matrix dimensions are as follows: $\tilde{A}$ has dimensions $(n\times n)$, $X^{(l)}$ is of dimensions $(n\times d)$, and $\mathcal{W}^{(l)}$ has dimensions $(d\times d)$. Time complexity for one layer of GNN computation is $\mathcal{O}(n^{2}d+nd^{2})$. If we take $L$ layers, then the total time is $\mathcal{O}(Ln^{2}d+Lnd^{2})$. The space complexity is $\mathcal{O}(n^{2}+Lnd+Ld^{2})$. When we compute on a sparse graph, the time complexity is $\mathcal{O}(m+Lnd^{2})$ and space complexity is $\mathcal{O}(m+Lnd+Ld^{2})$. The complexity is based on GCN. For other architectures like GAT veličković2018graphattentionnetworks, the complexity changes.
@@
The inference time complexity of our model is $\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]$, where $\bar{n}_{i}=n_{i}+\phi_{i}$ and $\phi_{i}$ is the number of additional nodes appended in each subgraph. Table 1 compares different approaches.
@@
The inference time complexity is at most $\sum_{i=1}^{k}[(n_{i}+\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}]\leq n^{2}d+nd^{2}$ if $\mathbb{E}[n_{i}+\phi_{i}]\leq\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\frac{d}{2}$.
@@
$\sum_{i=1}^{k}[(n_{i}+\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}]=\sum_{i=1}^{k}[n_{i}^{2}d+\phi_{i}^{2}d+2n_{i}\phi_{i}d+n_{i}d^{2}+\phi_{i}d^{2}]$
$=nd^{2}+d\left[\sum^{k}_{i=1}n_{i}^{2}+\sum^{k}_{i=1}\phi_{i}^{2}+2\sum_{i=1}^{k}n_{i}\phi_{i}+d\sum_{i=1}^{k}\phi_{i}\right]$
Now, define $\mathbb{E}[n_{i}]:=\frac{1}{k}\sum_{i=1}^{k}n_{i}=\frac{n}{k}=\frac{1}{r}$ (since $k=nr$) and $\mathbb{E}[\phi_{i}]:=\frac{1}{k}\sum_{i=1}^{k}\phi_{i}$. We also define the variances and covariance,
$\text{Var}(n_{i}):=\frac{1}{k}\sum_{i=1}^{k}n_{i}^{2}-\mathbb{E}[n_{i}]^{2}$ $\qquad\text{Var}(\phi_{i}):=\frac{1}{k}\sum_{i=1}^{k}\phi_{i}^{2}-\mathbb{E}[\phi_{i}]^{2},$
$\text{Cov}(n_{i},\phi_{i})=\text{Cov}$ $:=\frac{1}{k}\sum_{i=1}^{k}\phi_{i}n_{i}-\mathbb{E}[\phi_{i}]\mathbb{E}[n_{i}]$
@@
$\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\left(\frac{d}{2}+\frac{1}{r}\right)\geq 0$
$\Rightarrow\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})\geq\left(\frac{d}{2}+\frac{1}{r}\right)^{2}$

### v5

@@ 1 Introduction
To mitigate the high computational cost associated with the inference phase, we decompose the original graph $G$ into a collection of subgraphs $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$. This transformation yields substantial improvements in inference efficiency, achieving up to a $\mathbf{100\times}$ speedup for single-node prediction. Moreover, it significantly reduces memory consumption by as much as $\mathbf{100\times}$ —thereby enabling scalable inference on large-scale graphs comprising millions of edges. We provide a theoretical characterization of the upper bound on the expected size of subgraphs under which our method’s time and space complexity remain superior to that of the baselines. Notably, we observe that conventional methods often fail to perform inference on large graphs due to memory constraints, whereas our approach remains effective and scalable in such scenarios.
@@ 2 Related Work
The aim of reducing computational costs has been approached through various methods. Research by Chiang et al. (2019) and Zeng et al. (2020) uses a subgraph sampling technique, where a small subgraph is sampled at each training iteration for GNN training. Other methods, including layer-wise sampling and mini-batch training, have been explored by Chen et al. (2018a; b); Cong et al. (2020); Zou et al. (2019). Additionally, Fahrbach et al. (2020) introduced a coarsening algorithm based on Schur complements to create node embeddings in large graphs, while Huang et al. (2021) [SGGC] applied GNNs to coarsened graphs, reducing nodes and edges to lower training time. Kumar et al. (2023); Joly and Keriven (2024); Kataria et al. (2024); Dickens et al. (2024) also proposed different coarsening algorithms for scalable training approaches. Jin et al. (2021) [GCOND], Wang et al. (2024) [GC-SNTK] presented algorithms to generate a synthetic graph that mimics the original graph. However, their approach generates model-specific synthetic graphs that require training on the whole dataset, which defeats the purpose of reducing computational burden. Gupta et al. (2025) [BONSAI] addressed this issue by looking into computational trees and training on just the relevant computational trees. However, these approaches do not decrease inference time since the inference phase takes the whole graph as input. Xue et al. (2023) shifted the training process to subgraph-level, using coarsening algorithms to split the graph into $k$-subgraphs and recommending a multi-device training approach. In the graph-level task, Jin et al. (2022) [DOSCOND] and Xu et al. (2023) [KIDD] tackled the graph classification problem following up on the approach of GCOND and GC-SNTK. However, the graph regression problem was not explored in this direction.
@@ 3 Preliminaries
Graph Notations: Given an undirected graph $G=(V,E,X,W)$, $V$ is the vertex set, $E$ is the edge set, $X\in\mathbb{R}^{n\times d}$ is the feature matrix, and $W$ is the edge weight matrix. Let $|V|=n$ be the number of nodes and $|E|=m$ be the number of edges. Let $A\in\mathbb{R}^{n\times n}$ be the adjacency matrix of $G$ where $A_{ij}$ is the edge weight between nodes $v_{i}$ and $v_{j}$, and $d_{i}$ is the degree of the node $v_{i}$. $D\in\mathbb{R}^{n\times n}$ is a diagonal degree matrix with $i$th diagonal element as $d_{i}$. We denote $\mathcal{N}_{j}(v_{i})$ as the $j$-hop neighbourhood of node $v_{i}\in V$. Graph Neural Network: Graph Neural Network is a neural network designed to work for graph-structured data. The representation of each node in the graph is updated recursively by aggregation and transforming the node representations of its neighbors. Kipf and Welling (2017) introduced Graph Convolutional Network (GCN) as follows:
@@
Graph Coarsening and Graph Partitioning: Loukas (2019) discussed multiple coarsening algorithms which create a coarsened graph $G^{\prime}=(V^{\prime},E^{\prime},X^{\prime},W^{\prime})$ from a given graph $G=(V,E,X,W)$. We refer the vertex set $V^{\prime}=\{v^{\prime}_{1},v^{\prime}_{2},\dots,v^{\prime}_{k}\}$ of $G^{\prime}$ as coarsened nodes. Given a coarsening ratio $r\in[0,1]$, we have $\lfloor{n\times r}\rfloor, kk\in\mathbb{Z}$. We can interpret it as creating $k$ disjoint partitions, $C_{1},C_{2},\dots,C_{k}$, from a graph of $n$ nodes. Mathematically, we create a partition matrix $P\in\{0,1\}^{n\times k}$, where $P_{ij}=1$ if and only if node $v_{i}\in C_{j}$. $X^{\prime}=P^{\top}X$ is the coarsened node representation. $A^{\prime}=P^{\top}AP$ is the adjacency matrix of $G^{\prime}$. Similarly, the corresponding degree matrix $D^{\prime}=P^{\top}DP$. SGGC (Huang et al., 2021) used normalized partition matrix $\mathcal{P}=PC^{-\frac{1}{2}}$, where $C$ is defined as a diagonal matrix with diagonal entries $C_{jj}=|C_{j}|,j=1,2,\dots,k$.
Along with $G^{\prime}$, we create a set of disjoint subgraphs $\mathcal{G}_{s}=\{G_{1},G_{2},$ $\dots,G_{k}\}$ corresponding to the partitions $C_{1},C_{2},\dots,C_{k}$. The number of nodes in each partition $C_{i}$ is denoted by $n_{i}$. Each subgraph $G_{i}$ is the induced subgraph of $G$ formed by the nodes in $C_{i}$. $A_{i}$ and $D_{i}$ are the adjacency and degree matrix of $G_{i}$ respectively.
@@ 4 Methodology
In SGGC, a graph $G$ is reduced to $G^{\prime}$ using a partition matrix $P$ generated from the coarsening algorithm. The labels for the coarsened graph are $Y^{\prime}=\arg\max(P^{\top}Y)$, where $Y$ is the label matrix of $G$ storing the one-hot encoding of the label of each node for the node classification task. Training is carried out on the coarsened graph $G^{\prime}$ using $Y^{\prime}$ as the label information. While this approach is novel, essential label information is lost. For instance, suppose a node $v^{\prime}\in G^{\prime}$ is a combination of $v_{i},v_{j},v_{k}\in G$ where $v_{i},v_{j},v_{k}$ are test nodes. Let us assume that the classes of $v_{i},v_{j},v_{k}$ are $0,0,2$ respectively. $\arg\max$ will take the majority label and assign class 0 to $v^{\prime}$. Hence, the model will be trained to predict only the class 0, leading to the discarding of the quantification of the model’s performance on predicting less represented nodes.
@@
  • Extra Nodes: Adding the $1$-hop neighbouring nodes, namely Extra Nodes, in each subgraph $G_{i}$ denoted by $\mathcal{E}_{G_{i}}$. Xue et al. (2023) introduced this approach to reduce information loss after partitioning. We add a unit weight edge if two nodes in $\mathcal{E}_{G_{i}}$ are connected in $G$. Node embedding of $v\in\mathcal{E}_{G_{i}}$ is $x_{v}\in X$.
@@
  • Cluster Nodes: Although adding $1$-hop neighboring nodes in the subgraph helps recover lost information after partition, for multi-layered GNN models, information loss will persist. Instead of adding the neighboring nodes, we will add Cluster Nodes, which represent the neighboring clusters, denoted by $\mathcal{C}_{G_{i}}$. Liu et al. (2024) introduced this method to add nodes in the graph instead of adding to separate subgraphs. They also added cross-cluster edges among these Cluster Nodes and proposed a subgraph sampling-based method on the modified graph. In our work, we add cross-cluster edges. For each node $t\in\mathcal{C}_{G_{i}}$, the node embedding $x_{t}=X^{\prime}_{t}$.
@@
  • From Equation 2 and 3, we can say that $\sum_{i}|\mathcal{E}_{G_{i}}|\geq\sum_{i}|\mathcal{C}_{G_{i}}|$. This is because, using Extra Nodes, we add all the neighboring nodes. However, using Cluster Nodes, we only add one representative node for each neighboring cluster. Hence, the time taken to train and infer these modified subgraphs with Cluster Nodes is at most that with Extra Nodes.
@@
3:Construct $\mathcal{G}_{s}=\{G_{1}(V_{1},E_{1},X_{1}),G_{2}(V_{2},E_{2},X_{2}),\dots,G_{k}(V_{k},E_{k},X_{k})\}$ using $P$;
@@
5:  Append $\mathcal{E}_{G_{i}}$ or $\mathcal{C}_{G_{i}}$ to each $G_{i}$ $\triangleright$ Based on method of appending nodes
@@
8:Construct label matrix $\{Y_{1},Y_{2},\dots,Y_{k}\}$ for $\mathcal{G}_{s}$ ;
@@
10:for each subgraph $G_{i}$ in $\mathcal{G}_{s}$ do
11:  $O_{i}=M(L,A_{i},D_{i},X_{i})$; $\triangleright$ Algorithm 4
12:  $O.append(O_{i}[mask_{i}])$;
13:  $Y_{\textup{Collected}}.append(Y_{i}[mask_{i}])$;
14:end for
@@
16:Train $M$ to minimize LossLoss;
@@ 4.1 Node-level Task
We use Algorithm 4 with $L$ number of layers to create a standard GNN model $M$ for node-level tasks. While we use Algorithm 3 to train a model on the coarsened graph $G^{\prime}$, we propose Algorithm 1 to train a model on the set of subgraphs $\mathcal{G}_{s}$. Algorithm 3 utilizes the partition matrix $P$ generated by the coarsening algorithm to construct and train on the coarsened graph $G^{\prime}$ and the coarsened label information $Y^{\prime}$. On the other hand, our proposed Algorithm 1 uses the same partition matrix $P$ to construct and train on subgraphs $\mathcal{G}_{s}$ using original label information $Y$. For the node classification task, $\arg\max$ is used for creating $Y^{\prime}$ for $G^{\prime}$ and CrossEntropy as a loss function. For the node regression task, we do not create $G^{\prime}$. Mean Absolute Error (MAE) is used as the loss function.
@@ 4.2 Graph-level Task
For a given graph $G\in\mathcal{D}$, we create $G^{\prime}$ and $\mathcal{G}_{s}=\{G_{1},G_{2},$ $\dots,G_{k}\}$. Since $G^{\prime}$ is a single graph and $\mathcal{G}_{s}$ is a set of subgraphs, we use Algorithm 5 to create a model that trains on $G^{\prime}$ and Algorithm 2 to create a model that trains on $\mathcal{G}_{s}$. We use the CrossEntropy loss function for the graph classification task and the Mean Absolute Error (MAE) for the graph regression task.
@@
5:   $X_{i}^{(j)}=\sigma(\tilde{D_{i}}^{-\frac{1}{2}}\tilde{A_{i}}\tilde{D_{i}}^{-\frac{1}{2}}X_{i}^{(j-1)}\mathcal{W}^{(j-1)})$ $\triangleright$ Equation 1
@@
7:  $\tilde{X}.stack(X_{i}^{(L)})$ $\triangleright$ Stack along row axis
@@ 4.3 Time and Space Complexity
Train Inference
@@
[…]
@@
[…]
@@
[…]
@@
We can expand our inference time complexity as follows:
Now, define expected number of nodes in each subgraph $\mathbb{E}[n_{i}]:=\frac{1}{k}\sum_{i=1}^{k}n_{i}=\frac{n}{k}=\frac{1}{r}$ (since $k=nr$) and expected number of additional nodes appended in each subgraph $\mathbb{E}[\phi_{i}]:=\frac{1}{k}\sum_{i=1}^{k}\phi_{i}$. We also define the variances and covariance,
$\text{Var}(n_{i}):=\frac{1}{k}\sum_{i=1}^{k}n_{i}^{2}-\mathbb{E}[n_{i}]^{2},\quad\text{Var}(\phi_{i}):=\frac{1}{k}\sum_{i=1}^{k}\phi_{i}^{2}-\mathbb{E}[\phi_{i}]^{2},\quad\text{Cov}(n_{i},\phi_{i})=\text{Cov}:=\frac{1}{k}\sum_{i=1}^{k}\phi_{i}n_{i}-\mathbb{E}[\phi_{i}]\mathbb{E}[n_{i}]$
@@
$\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\left(\frac{d}{2}+\frac{1}{r}\right)\geq 0\Rightarrow\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})\geq\left(\frac{d}{2}+\frac{1}{r}\right)^{2}$